# Attribute Inference Attacks for Federated Regression Tasks

### v1

@@ B.2 Toy dataset setting
B.3 Data splitting strategy for Income-L dataset
@@ B.4 Hyperparameters
In our attack with an active adversary, detailed in Alg. 4, Adams’ hyperparameters are selected following a tuning process performed using Optuna (Akiba et al. 2019) hyperparameter optimization framework. In all the experiments, values for the learning rate, $\beta_{1}$ and $\beta_{2}$ are optimized to minimize each client’s training loss. Learning rate is tuned in range $[0.1,50]$, whereas $\beta_{1}$ and $\beta_{2}$ values are tested in range $[0.6,0.999]$. The optimal hyperparameters’ set is determined after 50 trials.
@@ C.1 Federated least squares regression
Here, we show the results of AIAs for federated least squares regression task, on Income-L, Income-A and Medical datasets. In all the experiments, each client trains a linear model for 300 communication rounds with 1 local epoch and batch size 32. The learning rate is set to $5\cdot 10^{-3}$. The passive adversary may eavesdrop all the exchanged messages until the end of the training. The active adversary launches the attack after 300 rounds for additional 10 and 50 rounds. All attacks are targeted at a randomly chosen single client. To approximate the optimal local model of the targeted client, our passive adversary applies Alg. 3 and our active adversary applies Alg. 4.
@@
We optimize the set of messages for both our method and gradient-based passive attacks. For gradient-based attacks, the set of inspected messages $\mathcal{T}$ is selected in $T^{p}=\{\{\text{first t}$ rounds in $\mathcal{T}_{c}\}$, for $t\in\{1,5,10,20,50,100,150,300\}\}$. For our attack, we explore a much larger number of sets of messages as our attack is computationally-light, evolving only simple matrix computation. In particular, we randomly sample $10^{7}$ sets of size $d+1$ and choose the one which minimizes the condition number of $\Theta_{\text{out}}$, since an ill-conditioned matrix $\Theta_{\text{out}}$ leads to a high numerical inversion error as mentioned in App. B.2. All other hyperparameters are consistent with those used in the neural network experiments.
@@
Results in Table 2 show that our attacks outperform the baseline Grad in both passive and active scenarios on across all three datasets. Notably, our passive attack achieves improvements of over 4, 5 and 7 percentage points (p.p.) for Income-L, Income-A and Medical datasets, respectively. Even when the gradient-based method has access to an oracle, our passive attacks still achieves higher accuracy on Income-A and Income-L and comes very close on Medical dataset. More importantly, our passive attack reaches already to the performance expected from an adversary who knows the optimal local model, demonstrating the effectiveness of our passive approach in approximating the optimal local model (Alg. 3). When shifting to active attacks, the performance gain over Grad remains largely consistent after 50 active rounds.

### v2

@@ B.3 Toy dataset setting
B.4 Data splitting strategy for Income-L dataset
@@ B.5 Hyperparameters
In our attack with an active adversary, detailed in Alg. 3, Adams’ hyperparameters are selected following a tuning process performed using Optuna (Akiba et al. 2019) hyperparameter optimization framework. In all the experiments, values for the learning rate, $\beta_{1}$ and $\beta_{2}$ are optimized to minimize each client’s training loss. Learning rate is tuned in range $[0.1,50]$, whereas $\beta_{1}$ and $\beta_{2}$ values are tested in range $[0.6,0.999]$. The optimal hyperparameters’ set is determined after 50 trials.
@@ C.1 Federated least squares regression
Here, we show the results of AIAs for federated least squares regression task, on Income-L, Income-A and Medical datasets. In all the experiments, each client trains a linear model for 300 communication rounds with 1 local epoch and batch size 32. The learning rate is set to $5\cdot 10^{-3}$. The passive adversary may eavesdrop all the exchanged messages until the end of the training. The active adversary launches the attack after 300 rounds for additional 10 and 50 rounds. All attacks are targeted at a randomly chosen single client. To approximate the optimal local model of the targeted client, our passive adversary applies Alg. 2 and our active adversary applies Alg. 3.
@@
We optimize the set of messages for both our method and gradient-based passive attacks. For gradient-based attacks, the set of inspected messages $\mathcal{T}$ is selected in $T^{p}=\{\{\text{first t}$ rounds in $\mathcal{T}_{c}\}$, for $t\in\{1,5,10,20,50,100,150,300\}\}$. For our attack, we explore a much larger number of sets of messages as our attack is computationally-light, evolving only simple matrix computation. In particular, we randomly sample $10^{7}$ sets of size $d+1$ and choose the one which minimizes the condition number of $\Theta_{\text{out}}$, since an ill-conditioned matrix $\Theta_{\text{out}}$ leads to a high numerical inversion error as mentioned in App. B.3. All other hyperparameters are consistent with those used in the neural network experiments.
@@
Results in Table 2 show that our attacks outperform the baseline Grad in both passive and active scenarios on across all three datasets. Notably, our passive attack achieves improvements of over 4, 5 and 7 percentage points (p.p.) for Income-L, Income-A and Medical datasets, respectively. Even when the gradient-based method has access to an oracle, our passive attacks still achieves higher accuracy on Income-A and Income-L and comes very close on Medical dataset. More importantly, our passive attack reaches already to the performance expected from an adversary who knows the optimal local model, demonstrating the effectiveness of our passive approach in approximating the optimal local model (Alg. 2). When shifting to active attacks, the performance gain over Grad remains largely consistent after 50 active rounds.